OpenStudy (anonymous):

Choose the correct simplification of the expression the (4x/y)^2. A. 4x^2/y^2 B. 8x^2/y^2 C. 16x^2/y^2 D. 16x^2/y

zepdrix (zepdrix):

When you have a square on the outside like this,\[\Large \left(\frac{ab}{c}\right)^2 \quad=\quad \frac{(ab)^2}{c^2} \quad=\quad \frac{a^2b^2}{c^2}\] You have to square eeeeeverything inside.